# Emergence of a hidden-order phase well below the charge density wave transition in a topological Weyl semimetal (TaSe4)2I

### v1

@@ I Introduction
The emergence and coexistence of distinct ordered states constitute a hallmark of strongly correlated systems, and deciphering their microscopic mechanisms lies at the forefront of condensed matter research. Symmetry breaking in such systems often drives the formation of new ordered states, manifested by the opening of an energy gap near the Fermi level ($E_{F}$) [gruner2018density, tinkham1996superconductivity, lee1973conductivity]. This gap is typically proportional to the magnitude of the corresponding order parameter. One well-known example is the charge density wave (CDW) phases in low-dimensional materials, characterized by periodic modulation of conduction electrons accompanied by a corresponding lattice distortion [peierls1955quantum, gruner2018density, monceau2012electronic]. CDW states often coexist with other emergent quantum phases, including superconductivity [Superconducting_CDW, Superconducting_CDW_TaS2], and ferromagnetism [FerromagneticCDW]. Furthermore, CDW systems serve as a versatile ground for realizing intriguing correlated phenomena such as the quantum anomalous Hall effect [AHall_CDW, FQAHall_CDW_PRB], axionic electrodynamics [AxionGooth2019], valley-selective transport [Valley_Polarization], and critical slowing down near phase transitions [Dynamical_Slowing_DownPRL].
@@
The fundamental collective excitations of the CDW state, amplitudon and phason, are manifestations of spontaneous symmetry breaking in the ordered phase [gruner2018density, Gedik_Collective]. Recently, the dynamics of these collective modes have been extensively investigated using a variety of experimental techniques [Dynamical_Slowing_DownPRL, kim2023observation, Fahad.CollectiveMode.PRL]. Owing to the finite gap in its dispersion, the amplitude mode contributes negligibly at low temperatures, except in the vicinity of the transition temperatures. By contrast, the phason is expected to be a gapless Goldstone mode; however, in real materials, impurities and defects lift this degeneracy, opening a finite gap and endowing the phason with mass [kim2023observation]. The resulting low-energy excitations, particularly those associated with mass-acquired phasons and hidden CDW orders, are of fundamental interest [Fahad.CollectiveMode.PRL, demsar1999single, Gedik_Collective].
@@
The broad family of transition metal tetrachalcogenides that have the generic formula (MX4)nY (where M = Nb, Ta; X = S, Se; Y = halogens such as Br, I; and n = 2, 3, 10/3) serves as quasi-one-dimensional model systems (quasi-1D) to study the role of filling the metal band $d_{z^{2}}$ [gressier1984characterization, Monceau1982, Roucau_1984, PhysRevB.94.104113, MMondalnTSI.PRB, Balandin2024, BALANDIN202274Review] giving rise to unconventional ground states, such as 1D band insulator [1DbandinsulatorPhysRevLett.84.1272], type III Weyl phase [li2021typeIIIWeyl], which hosts giant dielectric response [GiantdielectricPhysRevLett.96.046402], CDW state with boundary modes [Zahid_Nat.Phy] and so on. The highly correlated response of the ordered collective CDW modes in these materials can also give rise to complex low-energy dynamics.
@@
Among them, (TaSe4)2I (TSI2 in short) is the only type III CDW Weyl semimetal having a body-centered tetragonal crystal structure (space group-$I422$) at room temperature and undergoes the CDW transition near $T_{CDW}\sim 263~K$. As a rare Weyl-CDW material, TSI2 has recently sparked interest due to its potential to realize a so-called dynamical axion insulator state [AxionGooth2019], a correlated topological phase in which the CDW phase mode becomes an analog of the predicted axial field in high-energy physics. However, this interpretation is currently the subject of active debate [MonceauAPL2022, Zahid_Nat.Phy]. On the other hand, a recent study of ultrafast X-ray scattering suggested that the CDW amplitude mode is more than merely electronic; it also involves lattice degrees of freedom[Fahad.CollectiveMode.PRL]. Furthermore, TSI2 has also been observed to be an exceptional CDW candidate for the achievement of a stable 1D atomic chain limit, amporphorization superconductivity, and photodetection [wei2024polarization, li2024ultrabroadband, an2020long, Balandin2024]. These investigations unequivocally show that the topology and the electronic and lattice degrees of freedom of the collective CDW modes are strongly correlated in this compound. Despite this, all previous investigations in TSI2 have focused mainly on the $T_{CDW}\sim 263~K$ regime, while the low-temperature investigations of these correlated CDW collective modes are still very scarce.
@@
In this regard, low-frequency resistance noise spectroscopy has been demonstrated to be a sensitive tool for identifying phase transitions [chadniPhysRevLett.102.025701, majhi2021diffused, Kalimuddin.PRL, KaziAminPRL, Ghosh2023]. Near the transition, one can typically expect a divergent growth of the noise exponent ($\alpha$), resistance fluctuations $\left(\frac{\langle\delta R^{2}\rangle}{\langle R^{2}\rangle}\right)$, and relaxation time ($\tau$), a phenomenon often referred to as critical slowing down [Dynamical_Slowing_DownPRL, Satyaki_CriticalPRL]. The appearance of critical slowing down near the phase transition is a much more general phenomenon, as observed in first-order transitions [Satyaki_CriticalPRL], second-order transitions [Kalimuddin.PRL], and also in dynamical systems [Dynamical_Slowing_DownPRL]. Therefore, its widespread occurrence makes it a hallmark of phase transitions in a wide range of correlated solid-state systems.
@@ A Crystal structure and electrical transport
TSI2 single crystals have been grown using the chemical vapor transport (CVT) method using Iodine ($I_{2}$) as transport agent [gressier1984characterization, AxionGooth2019]. The details of the growth process are also available in our earlier report [Kalimuddin.PRL]. The crystal structure as inferred from single crystal Laue diffraction at room temperature reveals that TSI2 crystallizes in a body-centered tetragonal structure (space group I422, no. 97) (see S1 of Supplemental Material (SM) [supply]). The schematic representation of the projected tetragonal crystal structure has been shown in Fig. 1(a). The unit cell of TSI2 consists of parallel chains of TaSe4 units where Se4 of TaSe4 forms two dimers aligned in a plane, forming a rectangle. Each Ta atom is sandwiched at the center between two adjacent Se4 planes, which is presented in Fig. 1(b). Consecutive rectangular planes of Se4 are skew rotated by $\sim 45^{\circ}$. Infinitely long, well-separated, parallel skew-rotated TaSe4 chains provide a quasi-1D character and substantial anisotropy in the electronic structure.
TSI2 is currently the only known material system in which a CDW transition emerges from an underlying Weyl semimetal phase, potentially leading to the realization of an axion insulator state. In this system, the CDW—whose collective sliding mode is associated with an axion—is proposed to couple Weyl nodes of opposite chirality located at different points in the momentum space, thus breaking chiral symmetry and inducing an insulating gap, as illustrated in Fig. 1(c). This partial opening of the gap in TSI2 is caused by imperfect Fermi surface nesting [Shi2021], which may give rise to a secondary order significantly lower than $T_{CDW}$. The temperature-dependent normalized log resistivity is plotted in Fig. 1(d)(left). The CDW transition temperature ($T_{CDW}$) can be identified from the maxima of $\frac{dlog(R)}{d(10^{3}/T)}$ vs ($1/T$) plot in Fig. 1(d)(right). Notably, an additional, albeit weaker, resistivity anomaly is observed near $T^{*}$ $\sim$ 100 K, well below $T_{CDW}$, suggesting the possibility of a secondary phase transition. Although conventional transport measurements such as resistivity reveal only subtle signatures near $T^{*}$, low-frequency $1/f$ noise spectroscopy offers enhanced sensitivity to dynamical fluctuations, providing a more discerning probe to identify and characterize this secondary phase transition. Therefore, in the following subsection, we will primarily focus on the resistance noise measurement around $T^{*}$.
@@ B Low-frequency resistance noise measurement
Figure 2(a) shows the time traces of the resistance fluctuations for a few representative temperatures. Each data point for fluctuations has been vertically shifted by 0.2 % for clarity. Fluctuations increase in magnitude near the $T^{*}$ $\sim$100 K. The power spectral density (PSD) of the digitally filtered fluctuations within a bandwidth of 1 mHz–1 Hz is numerically calculated using Equation 1 employing Welch’s Periodogram method [ghosh2004set, Kalimuddin.PRL].
Figure 3: (a)-(c) log $\left(P(\delta R)\right)$ vs $\delta R^{2}$ for temperatures $^{*}, TTT\sim T^{*}$ and $T>T^{*}$ respectively. The solid (red) line is a guide to the Gaussian limit of the probability distribution of resistance fluctuations. The stochastic distributions of the fluctuations show extreme deviations from the normal distributions in the close vicinity of $T^{*}$. (d) A contour plot of PDF vs $\delta R$ with temperature. (e)-(g) Plot of normalized second spectra $s^{(2)}(f_{2})$ vs frequency for temperatures $^{*}, TTT\sim T^{*}$, and $T>T^{*}$ respectively. The solid (red) line represents the calculated Gaussian background of the second spectrum. (h) Plot of excess kurtosis ($\kappa=\sigma^{(2)}$-3) vs. temperature.
@@
The normalized PSD of the representative time series resistance fluctuations is plotted in Fig. 2(b). The noise magnitude and the spectral character show typical $1/f^{\alpha}$ behavior [CuerrentOpinion_Raychaudhuri, chadniPhysRevLett.102.025701, NbNPhysRevLett.111.197001], where $\alpha$ is the noise exponent. The temperature-dependent plot of $\alpha$ in Fig. 2(c) shows an increase upon approaching $T^{*}$ from $\alpha\sim 1.05$ at $T$ = 85 K to $\alpha\sim 1.45$ near 100 K. This rapid growth in $\alpha$ near $T^{*}$ suggests a strong shift of spectral weight to lower frequencies with a dominance of low-frequency/large-timescale of fluctuations [Satyaki_CriticalPRL, Kalimuddin.PRL]. The relative noise variance (integrated PSD) of the fluctuations $\left(\frac{\langle\Delta{R^{2}}\rangle}{\langle R^{2}\rangle}\right)$ within the band, fmin = 1 mHz and fmax = 1 Hz, is estimated using the relation 2[Raychaudhuri2002_basic, sudiptaPhysRevB.104.155101],
@@
The temperature variation of $\frac{\langle\Delta{R^{2}}\rangle}{\langle R^{2}\rangle}$ in Fig. 2(d) shows a rapid growth in the vicinity of $T^{*}$ by an order of magnitude compared to far temperature values. We have measured the low-frequency 1/f noise in another single crystal (sample II). Indeed, the signatures of the aforementioned hidden-order transition are reasonably reproduced in the second sample (see Fig. S2 of the SM[supply]). Remarkably, we noticed a similar behavior in the temperature-dependence of $S_{R}/R^{2}$, $\alpha$, and $\frac{\langle\Delta{R^{2}}\rangle}{\langle R^{2}\rangle}$.
The significant jump both in $\alpha$ and $\frac{\langle\Delta{R^{2}}\rangle}{\langle R^{2}\rangle}$ clearly indicate towards a phase transition near the $T^{*}$ [RevModPhysWeissman, Satyaki_CriticalPRL, Kalimuddin.PRL]. It is also noteworthy that here, the $T^{*}$ is significantly low compared to $T_{CDW}$, suggesting the order parameter is nearly fully gapped with a long-range ordered state. Thus, the random distribution of fluctuators is not entirely valid. This indicates a possibility of correlation in the fluctuators.
In order to identify the nature of the resistance fluctuations near the $T^{*}$, we have calculated the probability density function (PDF) of the detrended fluctuations [NbNPhysRevLett.111.197001]. For a Gaussian fluctuation with zero mean, the PDF (P($\delta R$)) is given by Eq. 3,
@@
Here, $\sigma^{2}$ represents the variance. Consequently, log(PDF) vs. $\delta R^{2}$ is a straight line, while the experimental time traces with correlation break such a pathway. Away from T, the distributions in $\delta R$ are mostly linear [Fig. 3(a,c)], suggesting Gaussian characteristics in $\delta R$. However, close to $T^{*}$, fluctuations have strong deviations from linear distribution, indicating a non-Gaussian distribution of resistance fluctuations, as shown in Fig. 3(b). In addition, the PDF broadens (growth in $\sigma$), which is a direct consequence of non-Gaussian fluctuations in the resistance noise [Swastik_PhysRevLett.91.216603]as shown in 3(d).
The frequency-domain determination of Kurtosis/normalized “second variance” ($k/\sigma^{(2)}$) allows an accurate determination of non-Gaussianity. We, therefore, calculate the second spectrum $\left(\Sigma^{(2)}(f_{2})\right)$ from the higher-order statistics of the resistance fluctuations. $\Sigma^{(2)}(f_{2})$ is a Fourier transform of the four-point correlation function, defined as $\Sigma^{(2)}(f_{2})$ = $\int_{0}^{\infty}\langle\delta R^{2}(t)\delta R^{2}(t+\varsigma)\rangle\cos(2\pi f\varsigma)d\varsigma$ [seidler1996non, Swastik_PhysRevLett.91.216603, chadniPhysRevLett.102.025701]. The non-Gaussian component in fluctuation is identified from the non-white character in $\Sigma^{(2)}(f_{2})$. It is a measure of ”spectral wandering” or fluctuations of the power spectrum itself [NbNPhysRevLett.111.197001]. The normalized second spectrum is given by Eq. 4,
$s^{(2)}(f)=\frac{\Sigma^{(2)}(f)df}{\big|\int_{f_{L}}^{f_{H}}S_{R}(f)df\big|^{2}}$ (4)
@@
Figure 3(h) shows the temperature variation of $k$. Interestingly, $k$ exhibits a pronounced peak near the $T^{*}$. This unambiguously denotes correlated fluctuations close to the $T^{*}$, where the fluctuations are slow. However, away from the transition region, $(\sigma^{(2)}-3)$ approaches 0. This is due to the uncorrelated Gaussian nature of the fluctuations [seidler1996non]. This is a direct signature of a phase transition accompanied by spatial correlation near $T^{*}$. Similar features of $\sigma^{(2)}$ are also evidenced near glassy freezing of electrons across the metal-insulator transition [Swastik_PhysRevLett.91.216603], Berezinskii-Kosterlitz-Thouless Transition [NbNPhysRevLett.111.197001], athermal phase transition [chadniPhysRevLett.102.025701] and Peierls transition [Kalimuddin.PRL].
@@ C Thermopower measurement
To gain further insight into the emergence of the hidden order phase around $T^{*}$, we measured Seebeck (S) as a function of temperature, since it carries information about the scattering associated with soft phonons [kuo2005anomalousseebeck] and finite changes in the density of states (DOS) near the Fermi surface [GeometryNatMat, Sruthi2022]. The temperature-dependence of $S(T)$ is described by well known Mott’s formula[Mott_PRB_1969], $S=-\frac{\pi^{2}k_{B}^{2}T}{3e}\left.\frac{d\ln\sigma(E)}{dE}\right|_{E=E_{F}}$, where $\sigma(E)$ represents energy-dependent electrical conductivity. Here, $k_{B}$, $e$, and $E_{F}$ represent the Boltzmann constant, electronic charge, and the Fermi energy, respectively.
@@
The temperature-dependent Seebeck in (TaSe4)2I was limited down to 115 K only [Bilusic2000], which effectively captured the CDW transition and anisotropy well. Here, we extend the Seebeck coefficient measurement down to $\sim$80 K, as shown in Fig. S5(a) (for measurement methodology, see S3 of the SM[supply]). The first derivative of $S(T)$ is also depicted in Fig. S5(b). S exhibits a linear and metallic behavior above the CDW transition and a nearly inverse $T$ variation below $T_{CDW}$. The crossover from linear to inverse $T$ dependence is marked by $T_{CDW}$, as shown in the inset of Fig. S5(a). At low temperatures, much below $T_{CDW}$, (TaSe4)2I exhibits an anomalous drop in S, suggesting a new phase transition around $T^{*}\sim$100 K. A similar anomaly is also observed in the temperature variation of $S(T)$ measured in another single crystal (sample II) (see S4 of the SM[supply]). Here, $T^{*}$ is evidenced from $dS/dT$, as illustrated in the inset of Fig. S5(b). The temperature-dependence of the Seebeck coefficient in a complex, correlated system like TSI is expected to have contributions from both the diffusion of charge carriers and phonon drag[MoO3_Thermopower]. The temperature-dependent Seebeck coefficient $S(T)$ in TSI2 is modeled to account for these two primary mechanisms. This is particularly important below the CDW transition at $T_{CDW}$, where strong electron-phonon coupling makes phonon drag a significant contributor.
We therefore fit $S(T)$ over the entire range to identify the contributions of different processes. Above $T_{CDW}$, TSI2 is a Weyl semimetal; consequently, $S(T)$ is linear in $T$ and can be described using the uncorrelated degenerate Fermi gas model, $\left(S(T)=-\frac{\pi^{2}k_{B}^{2}}{eE_{F}}T\right)$[Mott_PRB_1969, ziman2001electrons]. Below $T_{CDW}$, the Weyl nodes become gapped due to finite lattice distortion along the TaSe4 chains[van2001structure]. Considering the narrow band gap of (TaSe4)2I below and band-crossing above TCDW, we take into account the combined conduction for both electrons and holes (i.e, $\sigma=\sigma_{c}$$\sigma_{v}$) and fit $S(T)=\frac{\Delta_{eff}}{T}+const$, where $\Delta_{eff}$ represents the gap [fritzsche1971general]. The value of the estimated gap from Seebeck measurement is $\Delta_{eff}\sim 84$ meV.
The fit to Seebeck data deviates largely when $T$ is very close to both $T_{CDW}$ and $T^{*}$. The CDWs are electron-phonon coupling-driven phenomena, therefore expected to have a large contribution of strong phonon drag in $S(T)$. The complete functional form of $S(T)$ is given by [bailyn1960transport, cohn1991thermoelectric],

### v2

# PHYSICAL REVIEW B 113, 045114 (2026)

@@ I Introduction
The emergence and coexistence of distinct ordered states constitute a hallmark of strongly correlated systems, and deciphering their microscopic mechanisms lies at the forefront of condensed matter research. Symmetry breaking in such systems often drives the formation of new ordered states, manifested by the opening of an energy gap near the Fermi level ($E_{F}$) [26, 64, 37]. This gap is typically proportional to the magnitude of the corresponding order parameter. One well-known example is the charge density wave (CDW) phases in low-dimensional materials, characterized by periodic modulation of conduction electrons accompanied by a corresponding lattice distortion [50, 26, 45]. CDW states often coexist with other emergent quantum phases, including superconductivity [67, 40], and ferromagnetism [19]. Furthermore, CDW systems serve as a versatile ground for realizing intriguing correlated phenomena such as the quantum anomalous Hall effect [55, 60], axionic electrodynamics [23], valley-selective transport [43], and critical slowing down near phase transitions [71].
@@
The fundamental collective excitations of the CDW state, amplitudon and phason, are manifestations of spontaneous symmetry breaking in the ordered phase [26, 5]. Recently, the dynamics of these collective modes have been extensively investigated using a variety of experimental techniques [71, 32, 47]. Owing to the finite gap in its dispersion, the amplitude mode contributes negligibly at low temperatures, except in the vicinity of the transition temperatures. By contrast, the phason is expected to be a gapless Goldstone mode; however, in real materials, impurities and defects lift this degeneracy, opening a finite gap and endowing the phason with mass [32]. The resulting low-energy excitations, particularly those associated with mass-acquired phasons and hidden CDW orders, are of fundamental interest [47, 16, 5].
@@
The broad family of transition metal tetrachalcogenides that have the generic formula (MX4)nY (where M = Nb, Ta; X = S, Se; Y = halogens such as Br, I; and n = 2, 3, 10/3) serves as quasi-one-dimensional model systems (quasi-1D) to study the role of filling the metal band $d_{z^{2}}$ [24, 44, 53, 17, 9, 62, 8] giving rise to unconventional ground states, such as 1D band insulator [66], type III Weyl phase [39], which hosts giant dielectric response [61], CDW state with boundary modes [41] and so on. The highly correlated response of the ordered collective CDW modes in these materials can also give rise to complex low-energy dynamics.
@@
Among them, (TaSe4)2I (TSI2 in short) is the only type III CDW Weyl semimetal having a body-centered tetragonal crystal structure (space group-$I422$) at room temperature and undergoes the CDW transition near $T_{CDW}\sim 263~K$. As a rare Weyl-CDW material, TSI2 has recently sparked interest due to its potential to realize a so-called dynamical axion insulator state [23], a correlated topological phase in which the CDW phase mode becomes an analog of the predicted axial field in high-energy physics. However, this interpretation is currently the subject of active debate [59, 41]. On the other hand, a recent study of ultrafast X-ray scattering suggested that the CDW amplitude mode is more than merely electronic; it also involves lattice degrees of freedom[47]. Furthermore, TSI2 has also been observed to be an exceptional CDW candidate for the achievement of a stable 1D atomic chain limit, amporphorization superconductivity, and photodetection [68, 38, 2, 62]. These investigations unequivocally show that the topology and the electronic and lattice degrees of freedom of the collective CDW modes are strongly correlated in this compound. Despite this, all previous investigations in TSI2 have focused mainly on the $T_{CDW}\sim 263~K$ regime, while the low-temperature investigations of these correlated CDW collective modes are still very scarce.
@@
In this regard, low-frequency resistance noise spectroscopy has been demonstrated to be a sensitive tool for identifying phase transitions [12, 42, 30, 1, 22]. Near the transition, one can typically expect a divergent growth of the noise exponent ($\alpha$), resistance fluctuations $\left(\frac{\langle\delta R^{2}\rangle}{\langle R^{2}\rangle}\right)$, and relaxation time ($\tau$), a phenomenon often referred to as critical slowing down [71, 35]. The appearance of critical slowing down near the phase transition is a much more general phenomenon, as observed in first-order transitions [35], second-order transitions [30], and also in dynamical systems [71]. Therefore, its widespread occurrence makes it a hallmark of phase transitions in a wide range of correlated solid-state systems.
@@ A Crystal structure and electrical transport
TSI2 single crystals have been grown using the chemical vapor transport (CVT) method using Iodine ($I_{2}$) as transport agent [24, 23]. The details of the growth process are also available in our earlier report [30]. The crystal structure as inferred from single crystal Laue diffraction at room temperature reveals that TSI2 crystallizes in a body-centered tetragonal structure (space group I422, no. 97) (see S1 of Supplemental Material (SM) [56]). The schematic representation of the projected tetragonal crystal structure has been shown in Fig. 1(a). The unit cell of TSI2 consists of parallel chains of TaSe4 units where Se4 of TaSe4 forms two dimers aligned in a plane, forming a rectangle. Each Ta atom is sandwiched at the center between two adjacent Se4 planes, which is presented in Fig. 1(b). Consecutive rectangular planes of Se4 are skew rotated by $\sim 45^{\circ}$. Infinitely long, well-separated, parallel skew-rotated TaSe4 chains provide a quasi-1D character and substantial anisotropy in the electronic structure.
[…]
@@ B Low-frequency resistance noise measurement
Figure 2(a) shows the time traces of the resistance fluctuations for a few representative temperatures. Each data point for fluctuations has been vertically shifted by 0.2 % for clarity. Fluctuations increase in magnitude near the $T^{*}$ $\sim$100 K. The power spectral density (PSD) of the digitally filtered fluctuations within a bandwidth of 1 mHz–1 Hz is numerically calculated using Equation 1 employing Welch’s Periodogram method [21, 30].
[…]
@@
The normalized PSD of the representative time series resistance fluctuations is plotted in Fig. 2(b). The noise magnitude and the spectral character show typical $1/f^{\alpha}$ behavior [52, 12, 34], where $\alpha$ is the noise exponent. The temperature-dependent plot of $\alpha$ in Fig. 2(c) shows an increase upon approaching $T^{*}$ from $\alpha\sim 1.05$ at $T$ = 85 K to $\alpha\sim 1.45$ near 100 K. This rapid growth in $\alpha$ near $T^{*}$ suggests a strong shift of spectral weight to lower frequencies with a dominance of low-frequency/large-timescale of fluctuations [35, 30]. The relative noise variance (integrated PSD) of the fluctuations $\left(\frac{\langle\Delta{R^{2}}\rangle}{\langle R^{2}\rangle}\right)$ within the band, fmin = 1 mHz and fmax = 1 Hz, is estimated using the relation 2[52, 13],
@@
The temperature variation of $\frac{\langle\Delta{R^{2}}\rangle}{\langle R^{2}\rangle}$ in Fig. 2(d) shows a rapid growth in the vicinity of $T^{*}$ by an order of magnitude compared to far temperature values. We have measured the low-frequency 1/f noise in another single crystal (sample II). Indeed, the signatures of the aforementioned hidden-order transition are reasonably reproduced in the second sample (see Fig. S2 of the SM[56]). Remarkably, we noticed a similar behavior in the temperature-dependence of $S_{R}/R^{2}$, $\alpha$, and $\frac{\langle\Delta{R^{2}}\rangle}{\langle R^{2}\rangle}$.
The significant jump both in $\alpha$ and $\frac{\langle\Delta{R^{2}}\rangle}{\langle R^{2}\rangle}$ clearly indicate towards a phase transition near the $T^{*}$ [69, 35, 30]. It is also noteworthy that here, the $T^{*}$ is significantly low compared to $T_{CDW}$, suggesting the order parameter is nearly fully gapped with a long-range ordered state. Thus, the random distribution of fluctuators is not entirely valid. This indicates a possibility of correlation in the fluctuators.
In order to identify the nature of the resistance fluctuations near the $T^{*}$, we have calculated the probability density function (PDF) of the detrended fluctuations [34]. For a Gaussian fluctuation with zero mean, the PDF (P($\delta R$)) is given by Eq. 3,
@@
Here, $\sigma^{2}$ represents the variance. Consequently, log10(PDF) vs. $\delta R^{2}$ is a straight line, while the experimental time traces with correlation break such a pathway. Away from T, the distributions in $\delta R$ are mostly linear [Fig. 3(a,c)], suggesting Gaussian characteristics in $\delta R$. However, close to $T^{*}$, fluctuations have strong deviations from linear distribution, indicating a non-Gaussian distribution of resistance fluctuations, as shown in Fig. 3(b). In addition, the PDF broadens (growth in $\sigma$), which is a direct consequence of non-Gaussian fluctuations in the resistance noise [31]as shown in 3(d).
The frequency-domain determination of Kurtosis/normalized “second variance” ($k/\sigma^{(2)}$) allows an accurate determination of non-Gaussianity. We, therefore, calculate the second spectrum $\left(\Sigma^{(2)}(f_{2})\right)$ from the higher-order statistics of the resistance fluctuations. $\Sigma^{(2)}(f_{2})$ is a Fourier transform of the four-point correlation function, defined as $\Sigma^{(2)}(f_{2})$ = $\int_{0}^{\infty}\langle\delta R^{2}(t)\delta R^{2}(t+\varsigma)\rangle\cos(2\pi f\varsigma)d\varsigma$ [57, 31, 12]. The non-Gaussian component in fluctuation is identified from the non-white character in $\Sigma^{(2)}(f_{2})$. It is a measure of ”spectral wandering” or fluctuations of the power spectrum itself [34]. The normalized second spectrum is given by Eq. 4,
$s^{(2)}(f)=\frac{\Sigma^{(2)}(f)}{\big|\int_{f_{L}}^{f_{H}}S_{R}(f)df\big|^{2}}$ (4)
@@
Figure 3(h) shows the temperature variation of $k$. Interestingly, $k$ exhibits a pronounced peak near the $T^{*}$. This unambiguously denotes correlated fluctuations close to the $T^{*}$, where the fluctuations are slow. However, away from the transition region, $(\sigma^{(2)}-3)$ approaches 0. This is due to the uncorrelated Gaussian nature of the fluctuations [57]. This is a direct signature of a phase transition accompanied by spatial correlation near $T^{*}$. Similar features of $\sigma^{(2)}$ are also evidenced near glassy freezing of electrons across the metal-insulator transition [31], Berezinskii-Kosterlitz-Thouless Transition [34], athermal phase transition [12] and Peierls transition [30].
@@ C Thermopower measurement
To gain further insight into the emergence of the hidden order phase around $T^{*}$, we measured Seebeck (S) as a function of temperature, since it carries information about the scattering associated with soft phonons [36] and finite changes in the density of states (DOS) near the Fermi surface [27, 54]. The temperature-dependence of $S(T)$ is described by well known Mott’s formula[15], $S=-\frac{\pi^{2}k_{B}^{2}T}{3e}\left.\frac{d\ln\sigma(E)}{dE}\right|_{E=E_{F}}$, where $\sigma(E)$ represents energy-dependent electrical conductivity. Here, $k_{B}$, $e$, and $E_{F}$ represent the Boltzmann constant, electronic charge, and the Fermi energy, respectively.
@@
The temperature-dependent Seebeck in (TaSe4)2I was limited down to 115 K only [10], which effectively captured the CDW transition and anisotropy well. Here, we extend the Seebeck coefficient measurement down to $\sim$80 K, as shown in Fig. S5(a) (for measurement methodology, see S3 of the SM[56]). The first derivative of $S(T)$ is also depicted in Fig. S5(b). S exhibits a linear and metallic behavior above the CDW transition and a nearly inverse $T$ variation below $T_{CDW}$. The crossover from linear to inverse $T$ dependence is marked by $T_{CDW}$, as shown in the inset of Fig. S5(a). At low temperatures, much below $T_{CDW}$, (TaSe4)2I exhibits an anomalous drop in S, suggesting a new phase transition around $T^{*}\sim$100 K. A similar anomaly is also observed in the temperature variation of $S(T)$ measured in another single crystal (sample II) (see S4 of the SM[56]). Here, $T^{*}$ is evidenced from $dS/dT$, as illustrated in the inset of Fig. S5(b). The temperature-dependence of the Seebeck coefficient in a complex, correlated system like TSI is expected to have contributions from both the diffusion of charge carriers and phonon drag[63]. The temperature-dependent Seebeck coefficient $S(T)$ in TSI2 is modeled to account for these two primary mechanisms. This is particularly important below the CDW transition at $T_{CDW}$, where strong electron-phonon coupling makes phonon drag a significant contributor.
We therefore fit $S(T)$ over the entire range to identify the contributions of different processes. Above $T_{CDW}$, TSI2 is a Weyl semimetal; consequently, $S(T)$ is linear in $T$ and can be described using the uncorrelated degenerate Fermi gas model, $\left(S(T)=-\frac{\pi^{2}k_{B}^{2}}{eE_{F}}T\right)$[15, 70]. Below $T_{CDW}$, the Weyl nodes become gapped due to finite lattice distortion along the TaSe4 chains[65]. Considering the narrow band gap of (TaSe4)2I below and band-crossing above TCDW, we take into account the combined conduction for both electrons and holes (i.e, $\sigma=\sigma_{c}$$\sigma_{v}$) and fit $S(T)=\frac{\Delta_{eff}}{T}+const$, where $\Delta_{eff}$ represents the gap [20]. The value of the estimated gap from Seebeck measurement is $\Delta_{eff}\sim 84$ meV.
The fit to Seebeck data deviates largely when $T$ is very close to both $T_{CDW}$ and $T^{*}$. The CDWs are electron-phonon coupling-driven phenomena, therefore expected to have a large contribution of strong phonon drag in $S(T)$. The complete functional form of $S(T)$ is given by [7, 14],